OpenStudy (anonymous):

what is the horizontal asymptotes of (7x)/(x^4+1)^(1/4)

OpenStudy (dumbcow):

\[\lim_{x \rightarrow \infty} \frac{7x}{\sqrt[4]{x^4 + 1}} = \frac{7x}{x} = 7\]

OpenStudy (anonymous):

thats wrong